If 150cm3 of steam at 100∘C is added to 250cm3 water at 40∘C, how much steam(in cm3) is condensed into water at 100∘C?
1
Expert's answer
2020-12-06T17:28:10-0500

Answer

Heat present in steam =msL

​=150×540cal

=81000cal

Heat required by water to raise temperature from 40° to 100°

="m_s S\Delta T"

= 250×1×(100−40)

"=15000cal"

So

steam(in cm3) is condensed into water at 100∘C is

=250+"\frac{150\times5}{27}=277.77cm^3"